Write and equation of the line containing the given point and perpendicular to the given line. (5,-4);9x+7y=4
Answer by benazir.sj@gmail.com(109)

9x+7y =4
first we have to write the equation in standard form
y=mx+b

9x+7y=4
7y =-9x+4
y= -9/7x+4/7
here slope m1=-9/7
since the slope is perpendicular
m1*m2 =-1
m2=-1/m1
since the required line is perpendicular to 9x+7y=4
m2 =-1/m1
=-1/(-9/7)
=7/9
hence the equation of the requires line is
y-y1 =m2(x-x1)
(x1,y1)=(5,-4) m2=7/9
y-(-4) =7/9(x-5)
y+4=7/9x+7*5/9
y=7/9x+35/9-4
y=7/9x-1/9
or
9y=7x-1